# A $2$-distance $(2\Delta+7)$-coloring of planar graphs

### v2

@@ 2.2. The case $\Delta=7$
Let $v$ be a $5(4)$-vertex, and denote by $f_{1},f_{2},\ldots,f_{5}$ the faces incident to $v$ such that $f_{i}=v_{i}vv_{i+1}$ for $i\in[4]$. Remark that if $d_{2}(v)\leq 20$, then we set $G^{\prime}=G-v+\{v_{4}v_{5}\}$, and so the graph $G^{\prime}$ remains proper with respect to $G$. By minimality, the graph $G^{\prime}$ has a 2-distance coloring $f$ with $21$ colors. Since $d_{2}(v)\leq 20$, we can color $v$ with an available color, a contradiction. We therefore assume that $d_{2}(v)\geq 21$.
$(a)$ If $v$ has two 5(5)555(5)5 ( 5 )-neighbours, then there exist at least three edges in $G[N(v)]$ such that each of them is contained in two $3$-faces, and so we have $d_{2}(v)\leq 20$, a contradiction.
$(b)$ Suppose that $v$ has both a 5(5)555(5)5 ( 5 )-neighbour and a $4$-neighbour. Then we have again $d_{2}(v)\leq 20$, a contradiction.
$(c)$ Assume that $v$ has a $7(6)$-neighbour as well as $v$ is adjacent to two $4$-vertices. Since $v$ is adjacent to a $7(6)$-vertex, at least one edge in $G[N(v)]$ is contained in two $3$-faces. Thus, we have $d_{2}(v)\leq 20$, a contradiction.
$(d)$ Let $m_{4}(v)=0$. By contradiction, assume that $v$ has at most one neighbour consisting of $6(5^{-})$- or $7$-vertices, i.e., all neighbours of $v$ but one consist of 6(6)666(6)6 ( 6 )- or $5^{-}$-vertices. Notice that if $v$ has a 6(6)666(6)6 ( 6 )-neighbour, then there exist two edges in $G[N(v)]$ such that each of them is contained in two $3$-faces. It can be easily observed that $v$ has no four 6(6)666(6)6 ( 6 )-neighbours. Consider all the other possibilities; $v$ has either three 6(6)666(6)6 ( 6 )-neighbours or two 6(6)666(6)6 ( 6 )-neighbours and one $5^{-}$-neighbour or one 6(6)666(6)6 ( 6 )-neighbour and two $5^{-}$-neighbours or four $5^{-}$-neighbours. In each case, we have $d_{2}(v)\leq 20$, a contradiction. Thus, $v$ has two neighbours different from $5^{-}$- and 6(6)666(6)6 ( 6 )-vertex.
$(e)$ Let $m_{4}(v)=1$. Assume that $v$ has three neighbours $x,y,z$ consisting of $5^{-}$- or 6(6)666(6)6 ( 6 )-vertices. Recall that if $v$ has a 6(6)666(6)6 ( 6 )-neighbour, then there exist two edges in $G[N(v)]$ such that each of them is contained in two $3$-faces. Consider the possibilities; all $x,y,z$ are $5^{-}$-vertices; two of $x,y,z$ are $5^{-}$-vertices and one is 6(6)666(6)6 ( 6 )-vertex; one of $x,y,z$ is a $5^{-}$-vertex and the others are 6(6)666(6)6 ( 6 )-vertices; all $x,y,z$ are 6(6)666(6)6 ( 6 )-vertices. In each case, we observe that $d_{2}(v)\leq 20$, a contradiction.
$(f)$ Let $m_{4}(v)=1$. Suppose that $v$ has a 7(7)777(7)7 ( 7 )-neighbour. This implies that there exist two edges in $G[N(v)]$ such that each of them is contained in two $3$-faces. In such a case, if $v$ has two neighbours consisting of $5^{-}$- or 6(6)666(6)6 ( 6 )-vertices, then we would have $d_{2}(v)\leq 20$, a contradiction. That is, $v$ has at least four $6^{+}$-neighbours different from 6(6)666(6)6 ( 6 )-vertex. ∎
@@
    $v$ cannot have two neighbours consisting of $5(4)$-, 5(5)555(5)5 ( 5 )- or 6(6)666(6)6 ( 6 )-vertices. In particular, $v$ has at least one $7$-neighbour.
@@
    If $v$ has four $6$-neighbours, then $v$ has also a $7(5^{-})$-neighbour.
@@
    If $n_{5}(v)=0$, and $v$ has both a 6(6)666(6)6 ( 6 )-neighbour and a $6(5^{-})$-neighbour, then $v$ has three $7$-neighbours; at least one of which is a $7(5^{-})$-vertex.
@@
    If $n_{5}(v)=1$, then $v$ has no 6(6)666(6)6 ( 6 )-neighbours, in particular, $v$ has at most two $6$-neighbours.
@@
    If $n_{5}(v)=1$, and $v$ has exactly one $6(5^{-})$-neighbour, then $v$ has a $7(5^{-})$-neighbour.
@@
    If $n_{5}(v)=1$, and $v$ has exactly two $6(5^{-})$-neighbours, then $v$ has two $7(5^{-})$-neighbours.
@@
Let $v$ be a 5(5)555(5)5 ( 5 )-vertex, and denote by $f_{1},f_{2},\ldots,f_{5}$ the $3$-faces incident to $v$ such that $f_{i}=v_{i}vv_{i+1}$ for $i\in[5]$ in a cyclic fashion. Remark that if $d_{2}(v)\leq 20$, then we set $G^{\prime}=G-v$, and so the graph $G^{\prime}$ remains proper with respect to $G$. By minimality, the graph $G^{\prime}$ has a 2-distance coloring $f$ with $21$ colors. Since $d_{2}(v)\leq 20$, we can color $v$ with an available color, a contradiction. We therefore assume that $d_{2}(v)\geq 21$.
$(a)$ By Lemma 2.19(a), $v$ has no $4^{-}$-neighbours. Moreover, if $v$ has three $5$-neighbours, then we would have $d_{2}(v)\leq 19$, a contradiction.
$(b)$ If $v$ has two neighbours consisting of $5(4)$-, 5(5)555(5)5 ( 5 )- or 6(6)666(6)6 ( 6 )-vertices, then $d_{2}(v)\leq 20$, a contradiction. Besides, if all neighbours of $v$ are $6^{-}$-vertices, then we would have again $d_{2}(v)\leq 20$, a contradiction.
$(c)$ Suppose that $v$ has four $6$-neighbours. By $(b)$, the last neighbour of $v$ must be $7$-vertex, say $v_{1}$. If $v_{1}$ is a $7(6^{+})$-vertex, then there exists at least one edge in $G[N(v)]$ such that it is contained in two $3$-faces, and so we have $d_{2}(v)\leq 20$, a contradiction. Thus, $v_{1}$ is a $7(5^{-})$-vertex.
$(d)$ Let $n_{5}(v)=0$, and suppose that $v$ has both a 6(6)666(6)6 ( 6 )-neighbour and a $6(5^{-})$-neighbour. If $v$ has at most two $7$-neighbours, then we would have $d_{2}(v)\leq 20$, a contradiction. Moreover, at least one of those three $7$-neighbours must be $7(5^{-})$, since otherwise there exist at least three edges in $G[N(v)]$ such that each of them is contained in two $3$-faces, and so we have $d_{2}(v)\leq 20$, a contradiction.
$(e)$ Let $n_{5}(v)=1$. If $v$ has a 6(6)666(6)6 ( 6 )-neighbour or three $6$-neighbours, then we have $d_{2}(v)\leq 20$, a contradiction.
$(f)$ Let $n_{5}(v)=1$, and suppose that $v$ has exactly one $6(5^{-})$-neighbour. It follows from $(e)$ that $v$ has no 6(6)666(6)6 ( 6 )-neighbours, i.e., three neighbours of $v$ are $7$-vertices, where we recall that $v$ has no $4$-neighbour by $(a)$. If all those $7$-vertices are $7(6^{+})$-vertices, then there exist at least two edges in $G[N(v)]$ such that each of them is contained in two $3$-faces, and so we have $d_{2}(v)\leq 20$, a contradiction.
$(g)$ Let $n_{5}(v)=1$, and suppose that $v$ has exactly two $6(5^{-})$-neighbour, i.e., $v$ has two $7$-neighbours. If one of those $7$-vertices is a $7(6^{+})$-vertex, then there exists at least one edge in $G[N(v)]$ such that it is contained in two $3$-faces, and so we have $d_{2}(v)\leq 20$, a contradiction. Thus, $v$ has two $7(5^{-})$-neighbours.
$(h)$ Let $n_{5}(v)=2$. If $v$ has a $6$- or $7(6^{+})$-neighbours, then we have $d_{2}(v)\leq 20$, a contradiction. ∎
@@
Let $v$ be a $6(5)$-vertex. Then, $v$ has at most four bad $5$-neighbours. In particular, if $m_{4}(v)=1$ and $v$ has four bad $5$-neighbour, then $v$ has two $7(5^{-})$-neighbours.
@@
Let $v$ be a $6(5)$-vertex, and denote by $f_{1},f_{2},\ldots,f_{6}$ the faces incident to $v$ such that $f_{i}=v_{i}vv_{i+1}$ for $i\in[5]$.
Assume for a contradiction that $v$ has five bad $5$-neighbours. This implies that there exist at least three edges in $G[N(v)]$ such that each of them is contained in two $3$-faces, and so we have $d_{2}(v)\leq 19$. Since $v$ has five bad $5$-neighbours, one of $v_{1},v_{6}$ is a $5$-vertex, say $v_{1}$. Also, one of $v_{3},v_{5}$ is a $5$-vertex, say $v_{3}$. Let $G^{\prime}=G-v+\{v_{1}v_{6},v_{3}v_{1},v_{3}v_{5}\}$. Clearly, $G^{\prime}$ is proper with respect to $G$. By minimality, the graph $G^{\prime}$ has a 2-distance coloring $f$ with $21$ colors. Since $d_{2}(v)\leq 19$, we can color $v$ with an available color, a contradiction.
Now we suppose that $m_{4}(v)=1$, and $v$ has four bad $5$-neighbours. Since $v$ has four $5(4^{+})$-neighbours, there exist at least two edges in $G[N(v)]$ such that each of them is contained in two $3$-faces. In this manner, if $v$ has at most one $7(5^{-})$-neighbour, then we would have $d_{2}(v)\leq 20$. We get a contradiction as above. ∎
Let $v$ be a $7$-vertex and let $x$ be a $4$-neighbour of $v$. If $v$ and $x$ have two common neighbours, i.e., the edge $vx$ is contained in two $3$-faces, then $x$ is called a support vertex of $v$.
A vertex $v$ is said to be poor if it is a $4$- or 5(5)555(5)5 ( 5 )-vertex. Recall that $G$ has no 4(4)444(4)4 ( 4 )-vertices, and so poor vertices are $4(0)$-, $4(1)$-, $4(2)$-, $4(3)$- and 5(5)555(5)5 ( 5 )-vertices.
@@
    If $v$ has a support vertex, then $d_{2}(v)\geq 21$.
@@
    If $n_{4}(v)=3$, then $v$ has at most two 5(5)555(5)5 ( 5 )-neighbours.
@@
    If $n_{4}(v)=4$, then $v$ has at most one 5(5)555(5)5 ( 5 )-neighbours.
@@
    If $m_{3}(v)=5$, $n_{4}(v)=4$, $v$ has a support vertex, and $v$ has both a $5(4)$- and a 5(5)555(5)5 ( 5 )-neighbours, then $v$ is adjacent to a $7$-vertex.
@@
Let $v$ be a $7$-vertex, and denote by $f_{1},f_{2},\ldots,f_{7}$ the faces incident to $v$.
$(a)$ Suppose that $v_{1}$ is a support vertex of $v$. By definition, $v_{1}$ is a $4$-vertex and it is adjacent to both $v_{2}$ and $v_{7}$. Assume for a contradiction that $d_{2}(v)\leq 20$. When we set $G^{\prime}=G-v+\{v_{1}v_{3},v_{1}v_{4},v_{1}v_{5},v_{1}v_{6}\}$, the graph $G^{\prime}$ remains proper with respect to $G$. By minimality, the graph $G^{\prime}$ has a 2-distance coloring $f$ with $21$ colors. Since $d_{2}(v)\leq 20$, we can color $v$ with an available color, a contradiction.
$(b)$ Let $4\leq m_{3}(v)\leq 5$. Assume for a contradiction that all neighbours of $v$ are $4$-vertices, and so we have $d_{2}(v)\geq 20$. In addition, at least one neighbour of $v$ must be a support vertex due to $m_{3}(v)\geq 4$ and $n_{4}(v)=7$. It then follows from $(a)$ that we get a contradiction.
$(c)$ By contradiction, assume that $v$ has five poor neighbours. Note that a 5(5)555(5)5 ( 5 )-vertex is adjacent to at most one 5(5)555(5)5 ( 5 )-vertex by Lemma 2.24(b). Therefore $v$ has at most three 5(5)555(5)5 ( 5 )-neighbours as $m_{3}(v)=5$. Since $n_{4}(v)\leq 2$, and $v$ has five poor neighbours, we then deduce that $v$ has exactly three 5(5)555(5)5 ( 5 )-neighbours and two $4$-neighbours. However, this is not possible since a 5(5)555(5)5 ( 5 )-vertex has no $4$-neighbour by Lemma 2.24(a).
$(d)$ Let $v$ be a $7(6)$-vertex. By Lemma 2.24(a),(b), a 5(5)555(5)5 ( 5 )-vertex has no $4$-neighbour, also a 5(5)555(5)5 ( 5 )-vertex is adjacent to at most one bad $5$-vertex. Moreover, a $4(1^{+})$-vertex is adjacent to at most one $4$-vertex by Lemma 2.19(b). Those facts imply that $v$ cannot have six poor neighbours.
$(e)$ Let $n_{4}(v)=3$. By Lemma 2.24(a),(b), a 5(5)555(5)5 ( 5 )-vertex has no $4$-neighbour, also a 5(5)555(5)5 ( 5 )-vertex is adjacent to at most one bad $5$-vertex. Thus, one can infer that $v$ has at most two 5(5)555(5)5 ( 5 )-neighbours.
$(f)$ Similarly as $(d)$, the claim holds.
$(g)$ Suppose that $m_{3}(v)=5$, $n_{4}(v)=4$, $v$ has a support vertex, and $v$ has a $5(4)$- and 5(5)555(5)5 ( 5 )-neighbour. If the last neighbour of $v$ is a $6^{-}$-vertex, then we have $d_{2}(v)\leq 20$, a contradiction by $(a)$. ∎
We next apply discharging to show that $G$ does not exist. We use the same initial charges as before, with the following discharging rules.
@@
      $v$ gives $\frac{1}{5}$ to each of its $3$-neighbours,
@@
      $v$ gives $\frac{1}{4}$ to each of its bad $4$-neighbours,
@@
      $v$ gives $\frac{2}{9}$ to each of its 5(5)555(5)5 ( 5 )-neighbours,
@@
      $v$ gives $\frac{1}{9}$ to each of its $5(4)$-neighbours,
@@
      $v$ gives $\frac{1}{18}$ to each of its $6(5)$-neighbours.
@@
      $v$ gives $\frac{1}{6}$ to each of its 5(5)555(5)5 ( 5 )-neighbours,
@@
      $v$ gives $\frac{1}{9}$ to each of its $5(4)$-neighbours,
@@
      $v$ gives $\frac{1}{6}$ to each of its bad $4$-neighbours.
@@
      $v$ gives $\frac{1}{6}$ to each of its 5(5)555(5)5 ( 5 )-neighbours,
@@
      $v$ gives $\frac{1}{12}$ to each of its $5(4)$-neighbours.
@@
First we show that $\mu^{*}(f)\geq 0$ for each $f\in F(G)$. Given a face $f\in F(G)$, recall that the initial charge of a face $f$ is $\mu(f)=\ell(f)-4$. If $f$ is a $3$-face, then it receives $\frac{1}{3}$ from each of its incident vertices by R1, and so $\mu^{*}(f)\geq-1+3\times\frac{1}{3}=0$. If $f$ is a $4$-face, then $\mu(f)=\mu^{*}(f)=0$. Let $f$ be a $5^{+}$-face. By applying R2, $f$ sends $\frac{1}{5}$ to each of its incident vertices. It then follows that $\mu^{*}(f)\geq 1-5\times\frac{1}{5}=0$. Consequently, $\mu^{*}(f)\geq 0$ for each $f\in F(G)$.
We now pick a vertex $v\in V(G)$ with $d(v)=k$. By Lemma 2.16, we have $k\geq 3$.
(1). Let $k=3$. The initial charge of $v$ is $\mu(v)=d(v)-4=-1$. By Lemma 2.17, $m_{3}(v)=0$, $m_{4}(v)\leq 1$ and each neighbour of $v$ is a $7(5^{-})$-vertex. This means that $v$ is incident to at least two $5^{+}$-faces. It then follows that $v$ receives $\frac{1}{5}$ from each of its incident $5^{+}$-faces by R2, and at least $\frac{1}{5}$ from each of its $7(5^{-})$-neighbours by R6 and R7(a). Hence, $\mu^{*}(v)\geq-1+2\times\frac{1}{5}+3\times\frac{1}{5}=0$.

### v3

@@ 2.2. The case $\Delta=7$
Let $f_{i}=v_{i}vv_{i+1}$ for $i\in[4]$. If $d_{2}(v)\leq 20$, then setting $G^{\prime}=G-v+\{v_{4}v_{5}\}$ produces a graph $G^{\prime}$ that is proper with respect to $G$, contradicting Remark 2.1. Thus throughout the proof we assume $d_{2}(v)\geq 21$.
$(a)$ If $v$ had two $5(5)$-neighbours, then there would exist at least three edges in $E(v)$ such that each of them lies on two $3$-faces, which forces $d_{2}(v)\leq 20$ by Lemma 2.2, a contradiction.
$(b)$ If $v$ had both a $5(5)$-neighbour and a $4$-neighbour, then again $d_{2}(v)\leq 20$ by Lemma 2.2, a contradiction.
$(c)$ If $v$ had a $7(6)$-neighbour and two $4$-neighbours, then some edge of $E(v)$ would be contained in two $3$-faces, which gives $d_{2}(v)\leq 20$ by Lemma 2.2, a contradiction.
$(d)$ Let $m_{4}(v)=0$. Suppose that $v$ has at most one neighbour that is a $6(5^{-})$- or $7$-vertex. Then all remaining neighbours of $v$ are $6(6)$- or $5^{-}$-vertices. If $v$ has a $6(6)$-neighbour, then two edges of $E(v)$ lie in two $3$-faces. Considering all possible distributions of $5^{-}$- and $6(6)$-neighbours, one always obtains $d_{2}(v)\leq 20$ by Lemma 2.2, a contradiction. Thus $v$ must have at least two neighbours that are neither $5^{-}$- nor $6(6)$-vertices.
$(e)$ Let $m_{4}(v)=1$. Assume that $v$ has three neighbours forming $5^{-}$- or $6(6)$-vertices. Note that if $v$ has a $6(6)$-neighbour, then two edges of $E(v)$ lie on two $3$-faces. By checking all possible distributions of three neighbours among $5^{-}$-vertices and $6(6)$-vertices, we always obtain $d_{2}(v)\leq 20$ by Lemma 2.2, a contradiction.
$(f)$ Let $m_{4}(v)=1$ and suppose $v$ has a $7(7)$-neighbour. Then at least two edges of $E(v)$ are contained in two $3$-faces. If $v$ had two neighbours forming $5^{-}$- or $6(6)$-vertices, then $d_{2}(v)\leq 20$ by Lemma 2.2, a contradiction. ∎
@@
    $v$ cannot have two neighbours that are $5(4^{+})$-, or $6(6)$-vertices. In particular, $v$ has at least one $7$-neighbour.
@@
    If $v$ has four $6$-neighbours, then $v$ also has a $7(5^{-})$-neighbour.
@@
    If $n_{5}(v)=0$ and $v$ has both a $6(5^{-})$- and a $6(6)$-neighbour, then $v$ has three $7$-neighbours, at least one of which is $7(5^{-})$-vertex.
@@
    If $n_{5}(v)=1$, then $v$ has no $6(6)$-neighbours; in particular, $v$ has at most two $6$-neighbours.
@@
    If $n_{5}(v)=1$ and $v$ has exactly one $6(5^{-})$-neighbour, then $v$ has a $7(5^{-})$-neighbour.
@@
    If $n_{5}(v)=1$ and $v$ has exactly two $6(5^{-})$-neighbours, then $v$ has two $7(5^{-})$-neighbours.
@@
Let $f_{i}=v_{i}vv_{i+1}$ for $i\in[5]$ in cyclic order. If $d_{2}(v)\leq 20$, then taking $G^{\prime}=G-v$ gives a graph that is proper with respect to $G$, contradicting Remark 2.1. We therefore assume $d_{2}(v)\geq 21$.
$(a)$ By Lemma 2.21(a), $v$ has no $4$-neighbour. If $v$ had three $5$-neighbours, then $d_{2}(v)\leq 19$ by Lemma 2.2, a contradiction.
$(b)$ If $v$ had two neighbours consisting of $5(4^{+})$-, or $6(6)$-vertex, then $d_{2}(v)\leq 20$ by Lemma 2.2, a contradiction. Similarly, if all neighbours of $v$ were $6^{-}$-vertices, then again $d_{2}(v)\leq 20$, a contradiction. Hence $v$ must have at least one $7$-neighbour.
$(c)$ Suppose that $v$ has four $6$-neighbours. By (b), the fifth neighbour must be a $7$-vertex, say $v_{1}$. If $v_{1}$ were a $7(6^{+})$-vertex, then some edge in $E(v)$ would be contained in two $3$-faces, implying $d_{2}(v)\leq 20$ by Lemma 2.2, a contradiction. Hence $v_{1}$ is a $7(5^{-})$-vertex.
$(d)$ Let $n_{5}(v)=0$, and suppose that $v$ has both a $6(5^{-})$- and a $6(6)$-neighbour. If $v$ had only two $7$-neighbours, then $d_{2}(v)\leq 20$ by Lemma 2.2, a contradiction. Thus $v$ has three $7$-neighbours. At least one of them must be $7(5^{-})$-vertex; otherwise there would exist three edges of $E(v)$ such that each of them lies in two $3$-faces, again forcing $d_{2}(v)\leq 20$ by Lemma 2.2.
$(e)$ Let $n_{5}(v)=1$. If $v$ had a $6(6)$-neighbour or three $6$-neighbours, then $d_{2}(v)\leq 20$ by Lemma 2.2, a contradiction.
$(f)$ Let $n_{5}(v)=1$, and suppose that $v$ has exactly one $6(5^{-})$-neighbour. By (e), $v$ has no $6(6)$-neighbour, so the other three neighbours are $7$-vertices. If all were $7(6^{+})$-vertices, then at least two edges of $E(v)$ would lie in two $3$-faces, giving $d_{2}(v)\leq 20$ by Lemma 2.2, a contradiction. Thus $v$ has a $7(5^{-})$-neighbour.
$(g)$ Let $n_{5}(v)=1$, and suppose that $v$ has exactly two $6(5^{-})$-neighbours. By (a) and (e), $v$ has two $7$-neighbours. If one were a $7(6^{+})$-vertex, some edge of $E(v)$ would lie in two $3$-faces, giving $d_{2}(v)\leq 20$ by Lemma 2.2, a contradiction. Thus both $7$-neighbours are $7(5^{-})$-vertices.
$(h)$ Let $n_{5}(v)=2$. If $v$ had a $6$-neighbour or a $7(6^{+})$-neighbour, then $d_{2}(v)\leq 20$ by Lemma 2.2, a contradiction. ∎
@@
Let $v$ be a $6(5)$-vertex. Then $v$ has at most four bad $5$-neighbours. In particular, if $m_{4}(v)=1$ and $v$ has four bad $5$-neighbours, then $v$ has two $7(5^{-})$-neighbours.
@@
Let $f_{i}=v_{i}vv_{i+1}$ for $i\in[5]$. First, we claim that $v$ cannot have five bad $5$-neighbours. Indeed, if all five neighbours were bad $5$-vertices, then at least three edges in $E(v)$ would each lie in two $3$-faces, forcing $d_{2}(v)\leq 19$ by Lemma 2.2. Since $v_{1}$ or $v_{6}$ must be a $5$-vertex (say $v_{1}$) and similarly one of $v_{3},v_{5}$ must be a $5$-vertex (say $v_{3}$), we may take $G^{\prime}=G-v+\{v_{1}v_{6},\,v_{3}v_{1},\,v_{3}v_{5}\},$ and $G^{\prime}$ is proper with respect to $G$, contradicting Remark 2.1. Thus $v$ has at most four bad $5$-neighbours.
Now assume $m_{4}(v)=1$ and $v$ has four bad $5$-neighbours. Then at least two edges in $E(v)$ lie in two $3$-faces. If $v$ had at most one $7(5^{-})$-neighbour, then $d_{2}(v)\leq 20$ by Lemma 2.2, a contradiction. ∎
Let $v$ be a $7$-vertex, and let $x$ be a $4$-neighbour of $v$. If $v$ and $x$ have two common neighbours (equivalently, the edge $vx$ is contained in two $3$-faces), then $x$ is called a support neighbour of $v$.
A vertex is said to be poor if it is a $4$-vertex or a $5(5)$-vertex. Since $G$ has no $4(4)$-vertices, the poor vertices are exactly $4(0)$-, $4(1)$-, $4(2)$-, $4(3)$-, and $5(5)$-vertices.
@@
    If $v$ has a support neighbour, then $d_{2}(v)\geq 21$.
@@
    If $n_{4}(v)=3$, then $v$ has at most two $5(5)$-neighbours.
@@
    If $n_{4}(v)=4$, then $v$ has at most one $5(5)$-neighbour.
@@
    If $m_{3}(v)=5$, $n_{4}(v)=4$, $v$ has a support neighbour, and $v$ has both a $5(4)$- and a $5(5)$-neighbour, then $v$ must be adjacent to a $7$-vertex.
  • $(h)$
    If $m_{3}(v)=5$, $n_{4}(v)\geq 3$, $v$ has a support neighbour, and $v$ has a $3$-neighbour, then $v$ has at most four poor neighbours.
  • $(i)$
    If $m_{3}(v)=5$, $n_{4}(v)=5$, then $v$ has at most one $5(4)$-neighbour. In particular, if $v$ has a $5(4)$-neighbour or two $6(5)$-neighbours, then $v$ is incident to two $5^{+}$-faces.
  • $(j)$
    If $m_{3}(v)=5$, $n_{4}(v)=6$, then $v$ is incident to two $5^{+}$-faces; in particular, $v$ has no bad $5$-neighbour.
@@
$(a)$ Let $v_{1}$ be a support neighbour of $v$. Then $v_{1}$ is a $4$-vertex adjacent to both $v_{2}$ and $v_{7}$. Assume for a contradiction that $d_{2}(v)\leq 20$. Take $G^{\prime}=G-v+\{v_{1}v_{3},\,v_{1}v_{4},\,v_{1}v_{5},\,v_{1}v_{6}\}.$ Then $G^{\prime}$ is proper with respect to $G$, contradicting Remark 2.1. Thus $d_{2}(v)\geq 21$.
$(b)$ Let $4\leq m_{3}(v)\leq 5$. If all neighbours of $v$ were $4$-vertices, then $d_{2}(v)\geq 20$; and since $m_{3}(v)\geq 4$, at least one neighbour must be a support neighbour. By part (a), this yields a contradiction. Hence $v$ has at most six $4$-neighbours.
$(c)$ Assume $m_{3}(v)=5$ and $n_{4}(v)\leq 2$. Suppose for a contradiction that $v$ has five poor neighbours. By Lemma 2.26(b), a $5(5)$-vertex is adjacent to at most one $5(5)$-vertex. Since $m_{3}(v)=5$, $v$ can have at most three $5(5)$-neighbours. If $v$ has five poor neighbours, then it must have exactly three $5(5)$-neighbours and two $4$-neighbours, contradicting Lemma 2.26(a), which states that $5(5)$-vertices have no $4$-neighbours.
$(d)$ Let $v$ be a $7(6)$-vertex. By Lemma 2.26(a),(b), a $5(5)$-vertex has no $4$-neighbour and is adjacent to at most one bad $5$-vertex. Also, by Lemma 2.21(b), a $4(1^{+})$-vertex is adjacent to at most one $4$-vertex. These restrictions prevent $v$ from having six poor neighbours.
$(e)$ Let $n_{4}(v)=3$. Since a $5(5)$-vertex has no $4$-neighbour and is adjacent to at most one bad $5$-vertex (Lemma 2.26(a),(b)), $v$ can have at most two $5(5)$-neighbours.
$(f)$ Let $n_{4}(v)=3$. Then similar reasoning as in (d) implies that $v$ can have at most one $5(5)$-neighbour.
$(g)$ Assume $m_{3}(v)=5$, $n_{4}(v)=4$, $v$ has a support neighbour, and $v$ has both a $5(4)$- and a $5(5)$-neighbour. If the seventh neighbour of $v$ were a $6^{-}$-vertex, then $d_{2}(v)\leq 20$ by Lemma 2.2, contradicting part (a). Hence the remaining neighbour must be a $7$-vertex.
$(h)$ Assume the stated conditions. If $v$ had five poor neighbours, then $d_{2}(v)\leq 20$ by Lemma 2.2, contradicting (a).
$(i)$ Assume $m_{3}(v)=5$ and $n_{4}(v)=5$. Then $v$ has a support neighbour. If $v$ had two $5(4)$-neighbours, then $d_{2}(v)\leq 20$ by Lemma 2.2, contradicting (a). Thus $v$ has at most one such neighbour. Similarly, if $v$ has one $5(4)$-neighbour or two $6(5)$-neighbours, then $v$ must be incident to two $5^{+}$-faces.
$(j)$ Assume $m_{3}(v)=5$ and $n_{4}(v)=6$. Then $v$ has a support neighbour. If $v$ were not incident to two $5^{+}$-faces, then $d_{2}(v)\leq 20$ by Lemma 2.2, contradicting (a). Thus $v$ is incident to two $5^{+}$-faces. Similarly we deduce that $v$ has no $5(4)$- and $5(5)$-neighbours.. ∎
We now apply discharging to show that $G$ does not exist. We use the same initial charges as in the case $\Delta=6$, and redistribute charge according to the following rules.
@@
      $\frac{1}{5}$ to each of its $3$-neighbours,
@@
      $\frac{1}{4}$ to each of its bad $4$-neighbours,
@@
      $\frac{2}{9}$ to each of its $5(5)$-neighbours,
@@
      $\frac{1}{9}$ to each of its $5(4)$-neighbours,
@@
      $\frac{1}{18}$ to each of its $6(5)$-neighbours.
@@
      $\frac{1}{6}$ to each of its $5(5)$-neighbours,
@@
      $\frac{1}{9}$ to each of its $5(4)$-neighbours,
@@
      $\frac{1}{6}$ to each of its bad $4$-neighbours.
@@
      $\frac{1}{6}$ to each of its $5(5)$-neighbours,
@@
      $\frac{1}{12}$ to each of its $5(4)$-neighbours.
@@
First we show that $\mu^{*}(f)\geq 0$ for each $f\in F(G)$. Recall that every face $f$ has initial charge $\mu(f)=\ell(f)-4$. If $f$ is a $3$-face, then by R1 it receives $\frac{1}{3}$ from each of its incident vertices, and hence $\mu^{*}(f)\geq-1+3\times\frac{1}{3}=0$. If $f$ is a $4$-face, then $\mu(f)=0$ and it neither sends nor receives charge, so $\mu^{*}(f)=\mu(f)=0$. Let $f$ be a $5^{+}$-face. By R2, the face $f$ sends $\frac{1}{5}$ to each of its incident vertices. Therefore, $\mu^{*}(f)=\ell(f)-4-\ell(f)\times\frac{1}{5}=\frac{4\ell(f)}{5}-4\geq 0$. Consequently, every face $f\in F(G)$ satisfies $\mu^{*}(f)\geq 0$.
Next, let $v\in V(G)$ be a vertex of degree $d(v)=k$. By Lemma 2.18, we have $k\geq 3$.
(1). Let $k=3$. The initial charge of $v$ is $\mu(v)=d(v)-4=-1$. By Lemma 2.19, $m_{3}(v)=0$, $m_{4}(v)\leq 1$ and each neighbour of $v$ is a $7(5^{-})$-vertex. This implies that $v$ is incident to at least two $5^{+}$-faces. Then $v$ receives $\frac{1}{5}$ from each of its incident $5^{+}$-faces by R2, and at least $\frac{1}{5}$ from each of its $7(5^{-})$-neighbours by R6, R7(a). Hence, $\mu^{*}(v)\geq-1+2\times\frac{1}{5}+3\times\frac{1}{5}=0$.